# Exploring Time Delay Interferometry Ranging as a Practical Ranging Technique in the Bayesian Framework

### 05 Mar 2024 v1

@@ II.3 Parametrization of TDIR
TDIR utilizes parameterized functions as the hypothetical model, and the parameters will be constrained using the beatnote signals. Ref. [25] proposed a fitting form for the delays based on Keplerian orbit. This approach, according to our analysis, is only applicable when the delays are LTTs. Applying TDI to unsynchronized data demands that the delays also include information about clock deviations, hence a more general parametrization is desired. We employ a quadratic polynomial parametrization to fit the delays over a day:
@@
Figure 3: Left panel: the standard deviations of the differences between the fitting and true LTTs (orange curve) / pseudo ranges (green curve), evaluated per day in the whole mission period. Right panel: the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs. All the parameters are determined via least square fit.
The degree of polynomial is closely related to the time duration of data being processed. Due to the complexity of orbital dynamics and relativistic corrections, longer duration naturally requires higher-degree polynomials and more model parameters, hence heightening the challenge of MCMC parameter estimation. On the other hand, as will be analyzed in Sec. III, achieving better constraint on the delay parameters requires the availability of more data points. After balancing between difficulty and precision, we choose the duration of data to be one-day.
Through calculations based on the numerical orbit, we find that quadratic polynomial is sufficient for this timescale. The numerical orbit is provided by Ref. [44, 45], which takes into consideration the influence of the gravitational fields of celestial bodies and is qualified for space-based GW detection. Throughout the whole mission period of 5 years, we compute the LTTs using the formulas in Ref. [37], and fit Eq. (21) within each day using the least square fitting method. The standard deviations of the differences between the fitting and true LTTs are shown with the orange curve in the left panel of FIG. 3. Then, the original (uncorrected) clock deviations are set according to the USO model in Ref. [32], which assumes a $5\times 10^{-7}\ {\rm s/s}$ drift (43 ms/day). We repeat above calculations to the pseudo ranges resulting from this clock model, and the outcome is represented by the green curve in the same panel. For TDIR corresponding to the residual clock deviation left by ground-tracking synchronization, the result lies somewhere between the green and orange curves. According to the threshold established in Sec. II.2, we justify Eq. (21) to be a reasonable parametrization. The right panel of FIG. 3 shows the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs, with the parameters determined via least square fit. It is obviously shown that the result of fiducial TDIR manifests a very ideal LFN suppression effect. The residual LFN resulting from parametrized TDIR is well below 1/10 of the secondary noises, although larger than the fiducial one. The settings underlying this simulation are detailed in Sec. IV.1.
@@ III TDIR in the Bayesian framework
Figure 4: The ASDs of secondary noises and residual LFN in the $A_{2}$ and $T_{2}$ channels, with the armlengths calculated from the orbit data in the 1st day, and a 10 ns level ranging error. The shapes of $E_{2}$ channel are similar to those of $A_{2}$ hence not shown.
The data of a single TDI channel $X_{2}$ can only place constraints on $R_{12}$, $R_{21}$, $R_{13}$ and $R_{31}$. Using multiple channels is necessary for sampling the entire parameter space, and it can also help to obtain tighter constraints on the parameters. In further analysis based on the simulated data, to fully utilize the information from different TDI channels, and to maintain consistency with the conventions of GW source parameter estimation so as to be integrated into the global fit pipeline, we employ the “optimal” TDI channels constructed from the Michelson channels as [50]:
@@
and the total logarithmic likelihood function is the sum of $\{A_{2},E_{2},T_{2}\}$ channels. The PSDs of secondary noises in the optimal channels can be found in Appendix A, where we have also presented the PSDs of residual LFN due to the coupling with ranging errors. Observing Eq. (25), the contribution of a certain TDI channel to the total likelihood is determined by the ratio between residual LFN and secondary noises. These two terms for the $\{A_{2},T_{2}\}$ channels are shown in FIG. 4, with the armlengths calculated from the numerical orbit data in the 1st day, and a 10 ns level ranging error. The shapes of noises in $E_{2}$ are similar to those of $A_{2}$ hence not shown. As can be seen from the figure, within the frequency band in concern, each channel contributes similarly to the likelihood.
@@ IV.1 Settings for data simulation and MCMC sampling
Until now, the functional forms of residual clock errors left by ground-tracking synchronization for the Taiji mission are still under investigation. We assume their form as
@@
with the coefficients generated from normal distributions $\mathcal{C}_{m,i}\sim\mathcal{N}(0,0.1\ {\rm ms}/T_{\rm obs}^{m})$, $T_{\rm obs}=1\ {\rm day}$. Each term contributes 0.1 ms to the total deviation. According to the theoretical analysis in Sec. II, we do not expect the specific shape of $\delta\hat{\tau}_{i}^{t}(\tau)$ to have significant impact on the feasibility of TDIR. Regarding parametrization Eq. (21), the left panel of FIG. 3 shows that it is even applicable to the raw clock deviations, and the combination of different channels is valid up to the limits set at the end of Sec. III. Besides, we assume the ground-tracking range in each of the six laser links to deviate from corresponding LTT following the same form as Eq. (33). Note that these coefficients are independent of $\mathcal{C}_{m,i}$. These ranges are used as LTTs to calculate the theoretical PSDs in the likelihood. The simulation of unsynchronized beatnote signals in the frequency unit follows the approach described in Ref. [36, 38]. To be specific, the signals are initially generated in the proper time of each S/C, then transferred to TCB, and finally resampled according to the clock deviation model set by Eq. (33). To set a benchmark, we calculate the fiducial TDIR according to Eq. (4-5) and obtain the “true values” of parameters via least square fitting.
@@
We use the affine-invariant MCMC ensemble sampler emcee to draw samples from the likelihood [52]. The stretch parameter is set to the default value $a=2$. 72 walkers (4 times the number of parameters) are employed to fully explore the shape of posterior. The whole estimation process can be divided into two phases. The first phase is a preliminary search, where we downsample the data to 1 Hz, and the interpolation order is set to 7, all for the sake of efficiency. $f_{\rm max}$ is correspondingly reduced to 20 mHz to avoid interpolation error. The preliminary search starts from 1 ms intervals centered at the ground-tracking ranges. Once the preliminary search reaches a steady posterior distribution, the second phase, namely the refined search begins from the resultant $3\ \sigma$ confidence intervals. This phase is based on the original 4 Hz data, and we adopt the settings given in Sec. III, i.e. interpolation order 15 and $f_{\rm max}=0.5\ {\rm Hz}$. For the refined search, we run at least 50 auto-correlation times after burn-in following the instruction of the code.
@@ Appendix A Noise models
where $u_{i_{1}...i_{n}}\equiv 2\pi f(d_{i_{1}i_{2}}+...+d_{i_{n-1}i_{n}})$, $\Delta_{i_{1}...i_{n}}\equiv\exp(-iu_{i_{1}...i_{n}})$, $d_{i_{1}i2},...,d_{i_{n-1}i_{n}}$ being the LTTs. The expressions for $Y_{2}$, $Z_{2}$, $Y_{2}Z_{2}^{*}$ and $Z_{2}X_{2}^{*}$ can be obtained from Eq. (40) and Eq. (41) via cyclical permutation of the indices : $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$.

### 12 Mar 2024 v3

@@ II.3 Parametrization of TDIR
The TDIR algorithm utilizes parameterized function as the hypothetical model, and the parameters will be constrained using the beatnote signals. Ref. [25] proposed a fitting form for the delays based on Keplerian orbit. This approach, according to our analysis, is only applicable when the delays are LTTs. Applying TDI to unsynchronized data demands that the delays also include information about clock deviations, hence a more general parametrization is desired. We employ a quadratic polynomial parametrization to fit the delays over a day:
@@
Figure 3: Left panel: the standard deviations of the differences between the fitting and true LTTs (orange curve) / TDIRs (green curve), evaluated per day in the whole mission period. The TDIRs are calculated using on the original (uncorrected) clock deviations. Right panel: the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs. The simulation is conducted based on the 1st day of Taiji’s numerical orbit, and the uncorrected clock model. All the parameters are determined via least square fit.
The degree of polynomial is closely related to the time duration of data. Due to the complexity of orbital dynamics and relativistic corrections, longer duration naturally requires higher-degree polynomials and more model parameters, hence heightening the challenge of MCMC parameter estimation. On the other hand, as will be analyzed in Sec. III, achieving better constraint on the parameters requires the availability of more data points. After balancing between difficulty and precision, we choose the duration of data to be one day.
Through calculations based on the numerical orbit, we find that quadratic polynomial is sufficient for this timescale. The numerical orbit is provided by Ref. [44, 45], which takes into consideration the influence of the gravitational fields of celestial bodies and is qualified for space-based GW detection. Throughout the whole mission period of 5 years, we compute the LTTs using the formulas in Ref. [37], and fit Eq. (21) within each day using the least square fitting method. The standard deviations of the differences between the fitting and true LTTs are shown with the orange curve in the left panel of FIG. 3. Then, the original (uncorrected) clock deviations are set according to the USO model given in Ref. [32], which assumes a $5\times 10^{-7}\ {\rm s/s}$ drift (43 ms/day). We repeat above calculations to the fiducial TDIR resulting from this clock model, and the outcome is represented by the green curve in the same panel. For TDIR corresponding to the residual clock deviation left by ground-tracking synchronization, the result lies somewhere between the green and orange curves. According to the threshold established in Sec. II.2, we justify Eq. (21) to be a reasonable parametrization. The right panel of FIG. 3 shows the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs, with the parameters determined via least square fit. The simulation is conducted based on the 1st day of Taiji’s numerical orbit, and the uncorrected clock model. It is obviously shown that the result of fiducial TDIR manifests a very ideal effect of LFN suppression. The residual LFN resulting from parametrized TDIR is well below 1/10 of the secondary noises, although larger than the fiducial one.
@@ III TDIR in the Bayesian framework
Figure 4: The ASDs of secondary noises and residual LFN in the $A_{2}$ and $T_{2}$ channels, with the armlengths calculated from the orbit data in the 1st day, and the ranging error are set to 10 ns. The shapes of $E_{2}$ channel are similar to those of $A_{2}$ hence not shown.
Data of a single TDI channel $X_{2}$ can only place constraints on $R_{12}$, $R_{21}$, $R_{13}$ and $R_{31}$. Using multiple channels is necessary for sampling the entire parameter space, and it can also help to obtain tighter constraints on the parameters. In further analysis based on the simulated data, to fully utilize the information from different TDI channels, and to maintain consistency with the conventions of GW source parameter estimation so as to be integrated into the global fit pipeline, we employ the “optimal” TDI channels constructed from the Michelson channels as [50]:
@@
and the total logarithmic likelihood function is the sum of $\{A_{2},E_{2},T_{2}\}$ channels. The PSDs of secondary noises in the optimal channels can be found in Appendix A, where we have also presented the PSDs of residual LFN due to the coupling with ranging errors. Observing Eq. (25), the contribution of a certain TDI channel to the total likelihood is determined by the ratio between residual LFN and secondary noises. These two terms for the $\{A_{2},T_{2}\}$ channels are shown in FIG. 4, with the armlengths calculated from the numerical orbit data in the 1st day, and the ranging errors are set to 10 ns. The shapes of noises in $E_{2}$ are similar to those of $A_{2}$ hence not shown. As can be seen from the figure, within the frequency band in concern, each channel contributes similarly to the likelihood.
@@ IV.1 Settings for data simulation and MCMC sampling
Until now, the functional forms of residual clock errors left by ground-tracking synchronization for the Taiji mission are still under investigation. We treat the magnitudes reported in Ref. [22] as the state-of-the-art, and assume their form as
@@
with the coefficients generated from normal distributions $\mathcal{C}_{m,i}\sim\mathcal{N}(0,0.1\ {\rm ms}/T_{\rm obs}^{m})$, $T_{\rm obs}=1\ {\rm day}$. Each term contributes 0.1 ms to the total deviation. According to the theoretical analysis in Sec. II.1, we do not expect the specific shape of $\delta\hat{\tau}_{i}^{t}(\tau)$ to have a significant impact on the feasibility of TDIR. Regarding parametrization Eq. (21), the left panel of FIG. 3 shows that it is even applicable to the raw clock deviations, and the combination of different channels is valid up to the limits set at the end of Sec. III. Besides, we assume the ground-tracking range in each of the six laser links to deviate from corresponding LTT following the same form as Eq. (33). Note that these coefficients are independent of $\mathcal{C}_{m,i}$. These ranges are used as LTTs to calculate the theoretical PSDs in the likelihood. The simulation of unsynchronized beatnote signals in the frequency unit follows the approach described in Ref. [36, 38]. To be specific, the signals are initially generated in the proper time of each S/C, then transferred to TCB, and finally resampled according to the clock deviation model set by Eq. (33). To set a benchmark, we calculate the fiducial TDIR according to Eq. (4-5) and obtain the “true values” of parameters via least square fitting.
@@
We use the affine-invariant MCMC ensemble sampler emcee to draw samples from the likelihood [52]. The stretch parameter is set to the default value $a=2$. 72 walkers (4 times the number of parameters) are employed to fully explore the shape of posterior. The whole estimation process can be divided into two phases. The first phase is a preliminary search, where we downsample the data to 1 Hz, and the interpolation order is set to 7 for the sake of efficiency. $f_{\rm max}$ is correspondingly reduced to 20 mHz to avoid interpolation error. Following the evaluation in Sec. II.1, the preliminary search starts from 1 ms intervals centered at the ground-tracking ranges. Once the preliminary search reaches a steady posterior distribution, the second phase, namely the refined search begins from the resultant $3\ \sigma$ confidence intervals. This phase is based on the original 4 Hz data, and we adopt the settings given in Sec. III, i.e. interpolation order 15 and $f_{\rm max}=0.5\ {\rm Hz}$. For the refined search, we run at least 50 auto-correlation times after burn-in following the instruction of the code.
@@ Appendix A Noise models
where $u_{i_{1}...i_{n}}\equiv 2\pi f(d_{i_{1}i_{2}}+...+d_{i_{n-1}i_{n}})$, $d_{i_{1}i2},...,d_{i_{n-1}i_{n}}$ being the LTTs, and $\Delta_{i_{1}...i_{n}}\equiv\exp(-iu_{i_{1}...i_{n}})$. The expressions for $Y_{2}$, $Z_{2}$, $Y_{2}Z_{2}^{*}$ and $Z_{2}X_{2}^{*}$ can be obtained from Eq. (40) and Eq. (41) via cyclical permutation of the indices : $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$.